OpenStudy (anonymous):

Write the equation of a line in slope intercept form that passes through the points (0, 7) and (5, 3)

OpenStudy (welshfella):

do you know how to find the slope of the line joining these 2 points?

OpenStudy (madhu.mukherjee.946):

first find the slope slope=y2-y1/x2-x1 =3-7/5-0=-4/5

OpenStudy (madhu.mukherjee.946):

now use the formula= y-y1=m(x-x1) y-7=-4/5 (x-0) y-7=-4x/5 5y+4x=35

OpenStudy (madhu.mukherjee.946):

when two points are given slope is found by the formula